i) Find all points on the curve x(x + y2) = y where the tangent line is parallel to the x-axis.

ii) Find all points on the curve x(x + y2) = y where the tangent line is parallel to the y-axis


1
Expert's answer
2021-02-26T04:53:04-0500

i) Use implicit differentiation, then set dy dx = 0.


"\dfrac{d}{dx}(x(x+y^2))=\dfrac{dy}{dx}"

"2x+y^2+2xy\dfrac{dy}{dx}=\dfrac{dy}{dx}"

"\dfrac{dy}{dx}=\dfrac{2x+y^2}{1-2xy}"

"\dfrac{dy}{dx}=0=>\dfrac{2x+y^2}{1-2xy}=0"

"2x+y^2=0, 1-2xy\not=0"

"x=-\dfrac{y^2}{2}"

Substitute


"-\dfrac{y^2}{2}(-\dfrac{y^2}{2}+y^2)=y"

"-\dfrac{y^4}{4}=y"

"y_1=0, x_1=0, Point(0, 0)"


"y^3=-4"

"y_2=-\sqrt[3]{4}, x_2=-\sqrt[3]{2}, Point (-\sqrt[3]{2}, -\sqrt[3]{4})"


"Point(0, 0), Point (-\sqrt[3]{2}, -\sqrt[3]{4})"


ii) ) The points with vertical tangents are those where the denominator of "\dfrac{dy}{dx}" is zero (making the slope undefined). From part (i), we have


"\dfrac{dy}{dx}=\dfrac{2x+y^2}{1-2xy}=>1-2xy=0"

"x=\dfrac{1}{2y}, y\not=0"

Substitute


"\dfrac{1}{2y}(\dfrac{1}{2y}+y^2)=y"

"\dfrac{1}{4y^2}=\dfrac{y}{2}"

"y=\dfrac{\sqrt[3]{4}}{2}"

"x=\dfrac{\sqrt[3]{2}}{2}"

"Point(\dfrac{\sqrt[3]{2}}{2}, \dfrac{\sqrt[3]{4}}{2})"